OpenStudy (anonymous):

what is an angle between 0 and 2pi that is coterminal with 14pi/3

OpenStudy (anonymous):

You might wanna ask that on answers.yahoo XD but ill give it a try :3

OpenStudy (anonymous):

Recall that one can find angle coterminal to another angle by just adding or subtracting multiples of 2π. Since we want a coterminal angle that is positive and is between 0 and 2π, we can just add 2π to (-13/24)π to get the required coterminal angle to be: (-13/24)π + 2π = (-13/24 + 2)π = (35/24)π.